OpenStudy (anonymous):

If OB = 10 and RB = 8, then OR = ?

ganeshie8 (ganeshie8):

do u have a pic attached to this problem ?

ganeshie8 (ganeshie8):

u knw pythagoras theorem ?

OpenStudy (anonymous):

Umm sorta

ganeshie8 (ganeshie8):

good. we need to apply it here. the little square at R means, its a right angle. can u apply pythagoras theorem in triangle ORB ?

OpenStudy (anonymous):

I think so. So would it be a²+b²=c² ?

ganeshie8 (ganeshie8):

thats right. can u apply the same in triangle ORB ?

OpenStudy (anonymous):

Yes :) So I think It has something to do with 45-45-90, right?

OpenStudy (anonymous):

Yes :) So I think It has something to do with 45-45-90, right?

OpenStudy (anonymous):

Yes :) So I think It has something to do with 45-45-90, right?

ganeshie8 (ganeshie8):

its not that difficult. il work this out for u. see if it makes sense. in ORB right triangle, applying pythagoras theorem gives us : OB^2 = OR^2 + RB^2 10^2 = OR^2 + 8^2 100 = OR^2 + 64 OR^2 = 100-64 OR^2 = 36 OR = 6

OpenStudy (anonymous):

How'd you get 100-64?